# Revisting the Impact of Upstream Mergers with Downstream Complements and Substitutes

### v2

@@ Appendix A Proofs
I prove this lemma by providing counterexamples. Suppose that $n=3$, $c_{i}=0$ for $i=1,2,3$, and that demands are given as in equation (LABEL:eq:demandsEx) of the main text. When $i=1,2$, then $\partial D_{i}/\partial p_{j}=-b$ for $j\neq i$, whereas $\partial D_{3}/\partial p_{j}=-0.5\gamma/\sqrt{2-p_{1}-p_{2}}$ for $j=1,2$. Hence, if $b>0$ and $\gamma>0$, all three products are strict gross complements, while if $b<0$ and $\gamma<0$, all three products are strict gross substitutes.
@@
Thus, when $\gamma=1/2$, $\Delta=1.08+0.25-2\cdot(0.721)=-0.112<0$, whereas when $\gamma=1/2$, $\Delta=0.565+0.25-2\cdot(0.358)=0.099$. ∎
A.2 Proof of Proposition 3.2
@@
Fix $\bm{x}_{-12}$ and assume, without loss of generality, that among the products in $\bm{x}_{-12}$, the retailer does not carry the last $k\leq n-2$ products (note that $\bm{x}_{-12}$ has length $n-2$). If $k=n-2$, then the retailer only carries products 1 and 2. In this case, $\Pi^{*}(x_{1},x_{2},\bm{x}_{-12})$ exhibits strict increasing differences in $(x_{1},x_{2})$, following the exact same reasoning as in the proof of Proposition 3.1.
@@
Let $(q_{1}^{*},q_{2}^{*})$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ be the maximizers of $M(q_{1},q_{2})$, $M(q_{1},0)$, and $M(0,q_{2})$, respectively. By assumption, $q_{1}^{*}$, $q_{2}^{*}$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ are all strictly positive. Moreover, by construction, $\Pi^{*}(1,1,\bm{x}_{-12})=M(q_{1}^{*},q_{2}^{*})$, $\Pi^{*}(1,0,\bm{x}_{-12})=M(q_{1}^{\dagger},0)$, $\Pi^{*}(0,1,\bm{x}_{-12})=M(0,q_{2}^{\dagger})$, and $\Pi^{*}(0,0,\bm{x}_{-12})=M(0,0)$. Thus, $\Pi^{*}(x_{1},x_{2},\bm{x}_{-12})$ has strict increasing differences in $(x_{1},x_{2})$ if and only if:
@@
Appendix B An Example Illustrating Why There is No Analog to Proposition 3.2 for Gross Substitutes
In the main text, I claimed that it is not possible to derive a result analogous to Proposition 3.2 for the case of strict gross substitutes. This is because submodularity is not preserved under maximization. Here, I provide a concrete example to illustrate this phenomenon. Specifically, I demonstrate that even if products are strict gross substitutes and inverse demands are weakly submodular in $\bm{q}$, two products in the portfolio can still be complements for the retailer in terms of profits.
@@
Since $M(q_{1},q_{2})$ is strictly supermodular in $(q_{1},q_{2})$ this implies that when the retailer carries product 3, products 1 and 2 are strictly complements for the retailer in terms of profits (see the proof of Proposition 3.2). Indeed, in this particular case, $\Pi^{*}(1,1,1)+\Pi^{*}(0,0,1)-\Pi^{*}(1,0,1)-\Pi^{*}(0,1,1)=0.015>0$.
Finally, note that this issue does not arise when $\alpha>0$, $b>0$, and $\gamma>0$, i.e., when products are strict gross complements. In this case, $\sum_{i=1}^{i=3}P_{i}(\bm{q})q_{i}$ is always strictly supermodular in $\bm{q}$ and so is $\partial^{2}M/\partial q_{1}\partial q_{2}>0$. Indeed, as shown in Proposition 3.2, it is because $\sum_{i=1}^{i=3}P_{i}(\bm{q})q_{i}$ is strictly supermodular in $\bm{q}$ that $M(q_{1},q_{2})$ is strictly supermodular in $(q_{1},q_{2})$.

### v4

@@ A.1 Proof of Proposition 2
I prove this lemma by constructing examples. Suppose that $n=3$, $c_{i}=0$ for $i=1,2,3$, and that demands are given as in equation (7) of the main text. When $i=1,2$, then $\partial D_{i}/\partial p_{j}=-b$ for $j\neq i$, whereas $\partial D_{3}/\partial p_{j}=-0.5\gamma/\sqrt{2-p_{1}-p_{2}}$ for $j=1,2$. Hence, if $b>0$ and $\gamma>0$, all three products are strict gross complements, while if $b<0$ and $\gamma<0$, all three products are strict gross substitutes.
@@
Thus, when $\gamma=1/2$, $\Delta=1.08+0.25-2\cdot(0.721)=-0.112<0$, whereas when $\gamma=-1/2$, $\Delta=0.565+0.25-2\cdot(0.358)=0.099$. ∎
A.2 Proof of Proposition 3
@@
Fix $\bm{x}_{-12}$ and assume, without loss of generality, that among the products in $\bm{x}_{-12}$, the retailer does not carry the last $k\leq n-2$ products (note that $\bm{x}_{-12}$ has length $n-2$). If $k=n-2$, then the retailer only carries products 1 and 2. In this case, $\Pi^{*}(x_{1},x_{2},\bm{x}_{-12})$ exhibits strict increasing differences in $(x_{1},x_{2})$, following the exact same reasoning as in the proof of Proposition 2.
@@
Let $(q_{1}^{*},q_{2}^{*})$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ be the maximisers of $M(q_{1},q_{2})$, $M(q_{1},0)$, and $M(0,q_{2})$, respectively. By assumption, $q_{1}^{*}$, $q_{2}^{*}$, $q_{1}^{\dagger}$, and $q_{2}^{\dagger}$ are all strictly positive. Moreover, by construction, $\Pi^{*}(1,1,\bm{x}_{-12})=M(q_{1}^{*},q_{2}^{*})$, $\Pi^{*}(1,0,\bm{x}_{-12})=M(q_{1}^{\dagger},0)$, $\Pi^{*}(0,1,\bm{x}_{-12})=M(0,q_{2}^{\dagger})$, and $\Pi^{*}(0,0,\bm{x}_{-12})=M(0,0)$. Thus, $\Pi^{*}(x_{1},x_{2},\bm{x}_{-12})$ has strict increasing differences in $(x_{1},x_{2})$ if and only if:
@@
Appendix B An Example Illustrating Why There is No Analog to Proposition 3 for Gross Substitutes
In the main text, I claimed that it is not possible to derive a result analogous to Proposition 3 for the case of strict gross substitutes. This is because submodularity is not preserved under maximization. Here, I provide a concrete example to illustrate this phenomenon. Specifically, I demonstrate that even if products are strict gross substitutes and inverse demands are weakly submodular in $\bm{q}$, two products in the portfolio can still be complements for the retailer in terms of profits.
@@ Appendix B An Example Illustrating Why There is No Analog to Proposition 3 for Gross Substitutes
Since $M(q_{1},q_{2})$ is strictly supermodular in $(q_{1},q_{2})$ this implies that when the retailer carries product 3, products 1 and 2 are strictly complements for the retailer in terms of profits (see the proof of Proposition 3). Indeed, in this particular case, $\Pi^{*}(1,1,1)+\Pi^{*}(0,0,1)-\Pi^{*}(1,0,1)-\Pi^{*}(0,1,1)=0.015>0$.
Finally, note that this issue does not arise when $\alpha>0$, $b>0$, and $\gamma>0$, i.e., when products are strict gross complements. In this case, $\sum_{i=1}^{i=3}P_{i}(\bm{q})q_{i}$ is always strictly supermodular in $\bm{q}$ and so is $\partial^{2}M/\partial q_{1}\partial q_{2}>0$. Indeed, as shown in Proposition 3, it is because $\sum_{i=1}^{i=3}P_{i}(\bm{q})q_{i}$ is strictly supermodular in $\bm{q}$ that $M(q_{1},q_{2})$ is strictly supermodular in $(q_{1},q_{2})$.